A card is selected from a standard deck of 52 playing cards.Find probablilty of:
an even number ten or less given the card is not a heart
a five given the card is a face card
an Ace, given that the card is red

Answer by stanbon(75887)

A card is selected from a standard deck of 52 playing cards.
Find probablilty of:
an even number ten or less given the card is not a heart
# of ways to succeed: 3*5 = 15
P(conditions met) = 15/52
-----------------------------------
a five given the card is a face card
# of ways to succeed: 0
P(conditions met) = 0
-----------------------------------
an Ace, given that the card is red
# of ways to succeed: 2
P(conditions met) = 2/52 1/26
-------------------------
Cheers,
Stan H.